### v4

@@
Instrumented Difference-in-Differences with Heterogeneous Treatment Effectsthanks: This paper won the Kanematsu Prize from the Research Institution for Economics and Business Administration, Kobe University, in 2023. It also won the Best Young Scholar Award at SETA (Symposium on Econometrics and Applications) Conference, Taipei, in 2024. I am grateful to my advisors, Daiji Kawaguchi and Ryo Okui, for their continued guidance and support. I am also thankful to Alberto Abadie, Shakeeb Khan, Masamune Iwasawa, Shoya Ishimaru, Shosei Sakaguchi, Takahide Yanagi, and the three anonymous referees of the Kanematsu Prize for their insightful comments. I gratefully acknowledge the support provided by the JSPS KAKENHI Grant JP 24KJ0817. All errors are my own.
Sho Miyaji Graduate School of Economics, The University of Tokyo, 7-3-1 Hongo, Bunkyoku, Tokyo 113-0033, Japan; Email: shomiyaji-apple@g.ecc.u-tokyo.ac.jp.
(July 23, 2024)
Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). In contrast to Fuzzy DID design considered in de Chaisemartin and D’Haultfœuille (2018), our DID-IV design does not ex ante require strong restrictions on treatment adoption behavior across units. Additionally, our target parameter, the LATET, is policy-relevant if the instrument is based on the policy change of interest to the researcher. We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design, and propose an estimation method that is robust to treatment effect heterogeneity. We illustrate our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, which is the conventional approach to implement a staggered DID-IV design, yields a negative estimate, whereas our estimation method indicates a substantial gain from schooling.
@@ 1   Introduction
Instrumental variable (IV) strategies are widely used for causal inference across many fields in economics. Despite its popularity, empirical researchers commonly face two challenges in employing the IV method in practice. The first challenge lies in the internal validity of the IV estimate; finding an instrument that is effectively randomized across units, especially in observational studies, proves to be difficult. The second challenge lies in the external validity of the IV estimate; under heterogeneous treatment effects, a linear IV estimand captures the average treatment effect, for those who are affected by the instrument (Imbens and Angrist (1994)). However, this local average treatment effect (LATE) may not be policy-relevant if the instrument does not represent the policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
@@
To enhance at least the external validity of the IV estimate, many studies have exploited variation in the timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques. For example, Duflo (2001) estimates returns to schooling in Indonesia, leveraging variation arising from the new school construction program across regions and cohorts as an instrument for education attainment. As another example, Black et al. (2005) estimates the causal link between parents’ and children’s education attainment, using variation occurring from the different timing of implementation of school reform across municipalities and cohorts as an instrument for parents’ education attainment.
Notably, however, the underlying identification strategy behind these studies is inconsistent with IV designs: the instrument based on these policy changes is not randomized across units in reality, potentially threatening the internal validity of the IV estimate. Rather, the underlying identification strategy is similar to difference-in-differences (DID) designs, under where the effect of the policy shock (treatment) on the outcome is identified by exploiting the variation of that policy change across units and over time.
In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV). We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity in multiple time periods settings. Importantly, our DID-IV design does not require the randomization assumption of the instrument. In this design, we can achieve the internal and external validity of the IV estimate if the parallel trends assumptions in the treatment and the outcome are plausible in a given application, and the instrument is based on the policy shift of interest to the researcher.
@@
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument over two periods (unexposed group), whereas some units become exposed in the second period (exposed group). In this setting, our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period. Our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups. This framework comes from a simple and motivating observation: in the two-group/two-period ($2\times 2$) settings, a popular estimand, which we call the Wald-DID estimand, scales the DID estimand of the outcome by the DID estimand of the treatment between the two groups. We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
@@
In DID-IV designs, the interpretation of the parallel trends assumption in the outcome is less clear than that of the parallel trends assumption in DID designs because it does not depend on the untreated outcome. Rather, it depends on the outcome under no exposure to the instrument, which we call an unexposed outcome. Motivated by this observation, we then clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs. Specifically, we first uncover the selection mechanism behind DID-IV designs, showing that, in DID-IV settings, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of the policy shock. Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers. For instance, the time compliers do not adopt the treatment in the first period, but become adopt it in the second period without instrument. Subsequently, we show that the average time trend in the unexposed outcome equals a weighted average of the indirect effect of time on the outcome through treatment in each time group. We call this an expected time gain, and show that the parallel trends assumption in the outcome in DID-IV designs requires the same expected time gain between exposed and unexposed groups.
@@
de Chaisemartin and D’Haultfœuille (2018) (hereafter, ”dCDH”) also consider the same setting as in this article, but formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period. Accordingly, the identifying assumptions in dCDH are also different from those used here. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET under Fuzzy DID designs if the treatment effect is stable over time, and propose the alternative estimands that are free from the strong restriction on treatment effect heterogeneity.
@@
Next, motivated by the disagreement between this article and dCDH, we compare DID-IV designs to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs. First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units. These restrictions turn out to be asymmetric between exposed and unexposed groups, and difficult to defend from the institutional knowledge of the problem in a given application. Then, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers, who are affected by time but not affected by instrument. This decomposition result for the SLATET has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET in DID-IV designs. Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
We also revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and demonstrate that their argument is correct but misguided. First, we show that their argument stems from their reliance on the parallel trends assumption in DID designs. We then argue that this assumption is unsuitable for DID-IV settings for three reasons. First, this assumption generally fails to capture the average time trends of the outcome even in an unexposed group. This is because in DID-IV settings, units are allowed to receive the treatment in the absence of the policy shock (instrument). Second, in most applications of the DID-IV method, it is generally infeasible to impute the untreated potential outcomes (e.g., Duflo (2001), Black et al. (2005)). For instance, in Black et al. (2005), this assumption requires that the data contain the units with zero education attainment, which is unrealistic in practice. Finally, in DID-IV settings, we cannot directly test this assumption using the pre-exposed period data in general. Indeed, we show that the placebo test proposed by dCDH is incomplete for assessing the plausibility of this assumption.
@@
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In this setting, units start being exposed to the instrument at a particular point in time and remain exposed to that instrument afterward. Indeed, in most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), phased-in introduction of head starts across states (e.g. Johnson and Jackson (2019)), or gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions. Specifically, in this design, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument. Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings and state the interpretation of each assumption.
@@
We extend our DID-IV designs in several directions. We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment. We also consider extensions to repeated cross sections, and triple DID-IV designs. Lastly, we consider the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
@@
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable regressions. In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects under staggered DID-IV designs. Specifically, Miyaji (2024) shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects under staggered DID-IV designs, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$. We also discuss the procedure of pre-trend tests to check the validity of the parallel trends assumptions in the treatment and the outcome in DID-IV designs.
@@
We illustrate our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment. In this application, we first assess the plausibility of DID-IV identification assumptions implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. The TWFEIV regression performs the ”bad comparisons” (c.f. Goodman-Bacon (2021)) in the author’s setting: we treat the already exposed units as controls during the periods after both regions were affected by the policy change. Finally, we use our estimation method to reassess returns to schooling in the UK. We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighting estimate indicates roughly a $20\%$ gain from schooling in the UK.
@@
Overall, this paper provides a new econometric framework by combining the IV techniques with DID designs. Our DID-IV design can be applied to various empirical settings. First, it can cope with the general adoption process of the treatment when we are interested in the effect of the treatment on the outcome in DID designs. For instance, consider that some units already receive the treatment in period 00 before the policy shock, which distributes the treatment in period $1$. In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data. By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET. In other words, we can view that DID-IV is DID with the noncompliance of the assigned treatment.
@@
Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs. For instance, suppose that the treatment is education attainment and we are interested in returns to schooling as in Duflo (2001). In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality. However, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument). Once these groups are created under this design, one can identify the LATET by comparing the evolution of the treatment and the outcome between the two groups.
@@
Finally, our DID-IV design can overcome the endogeneity issue in empirical work. In practice, due to the endogenous adoption of the treatment, researchers often fail to identify the treatment effect via simple linear regressions or two-way fixed effects regressions. Using our DID-IV design allows addressing the common identification challenge by leveraging variation in the timing of policy adoption across units as an instrument for treatment. Importantly, in this design, the identifying variation comes not from the random assignment of the policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome. This identification strategy would be less demanding than IV designs in some applications, and enhances the internal and external validity of the IV estimate in practice.
@@
The rest of the paper is organized as follows. The following subsection discusses the related literature. Section 2 establishes DID-IV designs in a simple setting with two periods and two groups. Section 3 formalizes the target parameter and identifying assumptions in staggered DID-IV designs. Section 4 contains extensions. Section 5 presents our estimation method. Section 6 presents our empirical application. Section 7 concludes. All proofs are given in the Appendix.

### v5

@@
Abstract
Many studies exploit variation in policy adoption timing across units as an instrument for treatment. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. Moreover, we propose a credible estimation method in this design that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings, estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, the conventional approach to implement DID-IV designs, yields a negative estimate. By contrast, our estimation method indicates a substantial gain from schooling.
Abstract
[…]
@@ 1   Introduction
To identify the effect of a treatment on an outcome, many studies have exploited variation in policy adoption timing across units as an instrument for treatment. For example, to estimate returns to schooling in Indonesia, Duflo (2001) leverages variation in the timing of introduction of school construction program across regions as an instrument for education attainment. Similarly, to estimate the causal link between parents’ and children’s education attainment, Black et al. (2005) leverage variation in the timing of implementation of school reforms across municipalities as an instrument for parents’ education attainment. Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that the timing variation of a policy shock is used for identifying the treatment effects. Meanwhile, it also differs from DID designs, in that the variation is used to construct an instrument rather than a treatment.
In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV). We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity. We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this empirical application, we demonstrate that the estimation method matters in this design in practice.
@@
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument during the two periods (unexposed group), while some units become exposed in the second period (exposed group). [^1] Note that this two-period/two-group ($2\times 2$) setting is already considered in de Chaisemartin (2010) and Hudson et al. (2017), and our $2\times 2$ DID-IV design is mainly based on these studies. In this paper, we revisit $2\times 2$ DID-IV designs for two reasons. First, we aim to complement de Chaisemartin (2010) and Hudson et al. (2017) along two dimensions. Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs, and uncover the additional identifying assumption not give in the previous literature. In addition, we investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in DID-IV designs. Second, we aim to compare $2\times 2$ DID-IV designs to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018). While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) treat the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups. Our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period. We show that in this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the LATET.
@@
In DID-IV designs, the parallel trends assumption in the outcome requires careful assessment because it does not depend on the potential outcomes, but depends on the outcome under no exposure to the instrument, which we call an unexposed outcome. Give this, we next clarify the interpretation of this assumption in DID-IV settings. Specifically, we first point out that, in DID-IV designs, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of a policy shock (instrument). Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers. Subsequently, we show that the average time trend in an unexposed outcome equals a weighted average of the effect of time on an outcome through treatment in each time group. We call this an expected time gain, and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
@@
de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period. Because the target parameter differs from each other, the identifying assumptions in dCDH are also different from those used here.
@@
Motivated by the differences between this article and dCDH, we next examine the detailed connections between DID-IV and Fuzzy DID, and point out the issues inherent in Fuzzy DID. First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units, and these restrictions are asymmetric between unexposed and exposed groups. Second, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers. This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher. This is because the SLATET is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by instrument.
@@
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In most DID-IV applications, researchers exploit variation in policy adoption timing across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), phased-in introduction of head starts across states (e.g. Johnson and Jackson (2019)), or gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions. Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument. Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings, and state the interpretation of each assumption.
@@
We extend our DID-IV designs in several directions. We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment. We also consider extensions to repeated cross sections. Lastly, we consider the case where we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
@@
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable (TWFEIV) regressions (e.g. Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)). In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$. We also discuss the procedure of pre-trends tests to check the validity of the parallel trends assumptions in DID-IV designs.
@@
We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment. In this application, we first assess the plausibility of the DID-IV identification strategy implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. Finally, we use our estimation method to reassess returns to schooling in the UK. We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighting estimate indicates roughly a $20\%$ gain from schooling in the UK.
@@
Overall, this paper provides a new econometric framework by combining the IV techniques with DID designs. Our DID-IV design can be applied to various empirical settings. First, it can cope with the general adoption process of the treatment, which the canonical DID designs cannot, when we are interested in the effect of a treatment on an outcome. For instance, suppose that some units already receive the treatment in period 00 before the policy shock that distributes the treatment in period $1$. In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data. By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET. In other words, we can view that DID-IV is DID with noncompliance of the assigned treatment.
@@
Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs. For instance, suppose that the treatment is education attainment and we are interested in returns to schooling as in Duflo (2001). In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality. By contrast, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument). Once these groups are constructed, one can identify the local average treatment effect by comparing the evolution of the treatment and the outcome between the two groups.
@@
Finally, our DID-IV design allows a researcher to address the endogeneity issue in a given application. In many empirical settings, a researcher often fails to identify the treatment effects due to the endogenous adoption of the treatment or the potential violation of the parallel trends assumption under DID designs. In this situation, our DID-IV design offers an alternative identification strategy, constructing an instrument based on the timing variation of a policy shock. Notably, in this design, the identifying variation comes not from the random assignment of a policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome. This identification strategy would be less demanding than IV designs in certain applications, and enhance the internal and external validity of the IV estimate in practice (Imbens and Angrist (1994); Heckman and Vytlacil (2001); Heckman and Vytlacil (2005)).
@@
The rest of the paper is organized as follows. The following subsection discusses the related literature. Section 2 establishes DID-IV designs in two periods and two groups settings. Section 3 formalizes the target parameter and identifying assumptions in staggered DID-IV designs. Section 4 contains extensions. Section 5 presents our estimation method. Section 6 presents our empirical application. Section 7 concludes. All proofs are given in the Appendix.